Choose an integer, 4, between 85 and 95, and another integer, B, between 250 and 350. Determine the angle required for a baseball to leave the bat with an initial velocity of 4 miles per hour and first hit the ground B feet horizontally from home plate if contact is made 3 feet above the ground.


1
Expert's answer
2021-04-26T17:09:49-0400
"v=90\frac{5280}{3600}=132\frac{ft}{s}"

"300=132t\cos{\theta}\\t=\frac{300}{132\cos{\theta}}\\90=(132\sin{\theta})\frac{300}{132\cos{\theta}}-16\frac{300^2}{132^2\cos^2{\theta}}\\\theta=35.5\degree"